OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form

OpenStudy (anonymous):

y=2x+7

OpenStudy (precal):

y=mx+b m=2 x=-1 y=5 5=2(-1) +b 5=-2+b 7=b y=2x+7